OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of (f + g)(−2)? −9 −6 −3 9

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

is that X2 or X^2

OpenStudy (anonymous):

X^2

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

r u sure????? kenzmyers?

OpenStudy (anonymous):

if its right ill give the correct answer best response

OpenStudy (anonymous):

i am positive

OpenStudy (anonymous):

look

OpenStudy (anonymous):

can you help me with another question?

OpenStudy (anonymous):

kenzmyers?

OpenStudy (anonymous):

add the two functions , i.e. f(x) + g(x) then sub in -2 f+g = x^2 + 2x -9 f+g (-2) = 2(-2) + 1 +(-2)^2 -10 = -9

OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of (f + g)(−2)? so essentially you will evaluate each individual equation and then combine the two thus being: f: 2(-2) +1 = -3 g: (-2)^2 -10= -6 -3 + -6 =-9

OpenStudy (anonymous):

ohh ok thks kenzmyer and fortune

OpenStudy (anonymous):

will u help with another question?

OpenStudy (anonymous):

Choose the best description of how to calculate f[g(3)] if f(x) = 3x − 2 and g(x) = x2 − 1. Calculate f(3) and g(3), then multiply them together. Calculate f(3) and g(3), then write the results as an ordered pair. Calculate f(3) and g(3), then subtract them. Calculate g(3), plug the result into f(x) and simplify.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok i posted it on this

OpenStudy (anonymous):

Calculate g(3), plug the result into f(x) and simplify.

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of f[g(3)]? 39 −3 −7 −1

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i chose u as best response since ur helping me the most. thks so much

OpenStudy (anonymous):

first calculate g(3); g(3) = 3^2 - 10 = -1 then put that into f(x) = 2x + 1 f(g(3)) = 2(-1) + 1 = -1 final answer = -1

OpenStudy (anonymous):

I have one last question but Im going to post it as a new question so if you want to go see it and answer it ill chose u as best answer again.

OpenStudy (anonymous):

i became ur fan

OpenStudy (anonymous):

Lol thanks

OpenStudy (anonymous):

whats the question?